# Solving Inverse Problems in Stochastic Self-Organising Systems through Invariant Representations

### v1

@@ 5.1 Limitations
The embedding representations are only useful to the extent that they are capable of encoding the invariances present in the target patterns. If the embedding model fails to capture the relevant features of the patterns–or encode non-relevant invariances–the optimisation process may converge to parameters that reproduce visually dissimilar results, despite low embedding losses. For example, an embedding might prioritise shape over size, which can lead to false positives in parameter recovery. In other words, the embedding model is capturing an invariance, but might not be an invariance that we want. While this property may be desirable for certain problems, it can pose challenges for others. An example of this issue is illustrated in Fig. 9 where the technique finds a set of parameters that reproduce the target patterns but at a reduced scale.
Difficulty of quantitative evaluation.
The method comes with the challenge of quantitatively evaluating its results because evaluating the visual similarity between two patterns is fundamentally subjective [50]. Visual embedding distances provide a proxy for this, but they do not always match how humans perceive similarity. Often, judging success comes down to visual inspection–you know it when you see it. Although this works in practice, it is inherently subjective. Domain-specific metrics will be needed in order to make the evaluation more systematic.
@@ 5.2 Future Work
A future path is the use of custom embedding models trained specifically on the types of patterns relevant to the target domain. This could involve fine-tuning or training embedding models—e.g., a variational autoencoder trained on a domain-specific dataset [38]—or augmenting the dataset with relevant transformations for which we want the embedding model to be invariant. This way, the embedding space can be aligned more closely with the visual equivalence classes that matter for the task. Domain-specific embeddings could provide a way of ensuring that the method results in perceptually meaningful reconstructions.

### v2

@@ 5.1 Limitations
The embedding representations are only useful to the extent that they are capable of encoding the invariances present in the target patterns. If the embedding model fails to capture the relevant features of the patterns–or encode non-relevant invariances–the optimization process may converge to parameters that reproduce visually dissimilar results, despite low embedding losses. For example, an embedding might prioritise shape over size, which can lead to false positives in parameter recovery. In other words, the embedding model is capturing an invariance, but might not be an invariance that we want. While this property may be desirable for certain problems, it can pose challenges for others. An example of this issue is illustrated in Fig. 10 where the technique finds a set of parameters that reproduce the target patterns but at a reduced scale.
@@ 5.2 Future Work
One future direction is the use of custom embedding models trained specifically on the types of patterns relevant to the target domain. This could involve fine-tuning or training embedding models—e.g., a variational autoencoder trained on a domain-specific dataset [21]—or augmenting the dataset with relevant transformations for which we want the embedding model to be invariant. This way, the embedding space can be aligned more closely with the visual equivalence classes that matter for the task. Domain-specific embeddings could provide a way of ensuring that the method results in perceptually meaningful reconstructions.
@@
6 Acknowledgements
We thank Milton L. Montero for the discussion on the anisotropy of the reaction-diffusion model. This project was supported by the Novo Nordisk Foundation Synergy Grant REPROGRAM number NNF23OC0086722.